I am unsure if this is correct, please check and complete the net income/ explain how its done!! Thank you!!Junktrader is an online company that specializes in matching buyers and sellers of used items. Buyers and sellers can purchasRequired: 1. Prepare journal entries for the above transactions, which occurred during a recent month. 2. Determine Junktrade
Answer #1

Answer 1:

No Debit Credit Transaction a. General Journal Accounts Receivable Advertising (or Service) Revenue 365 365 80 Advertising Ex

Answer 2:

Preliminary Net Income $ 648

Working:

Net Income = Revenue - Expense

From above journals we find:

Revenue = 365 + 320 + 188 = $873

Expenses = 80 + 145 = $225

Net Income = 873 - 225 = $648